OpenStudy (anonymous):

Write Into a algebraic expression? Five times the difference of one-half of a number and two

OpenStudy (anonymous):

i guess we need a variable right?

OpenStudy (anonymous):

\(x\) is what the math teachers use so lets use that

OpenStudy (anonymous):

if that is the "number" then one half that number is \(\frac{x}{2}\)

OpenStudy (anonymous):

the difference between that number and 2 is \[\frac{x}{2}-2\]

OpenStudy (anonymous):

five times that is \[5\left(\frac{x}{2}-2\right)\]